Can someone help me solve this problem? Thank you so much!
Solve x² + 4x – 7 = 0 by completing the square. How many solutions are there?

Answer by scott8148(6628)

x^2 + 4x = 7

x^2 + 4x + (4/2)^2 = 7 + (4/2)^2

(x + 2)^2 = 11

x + 2 = ± √11

x = -2 ± √11